# Think Before Recommend: Unleashing the Latent Reasoning Power for Sequential Recommendation

### v1

@@ 1. Introduction
Recommender systems (RS) have become ubiquitous in modern daily life, powering personalized services across domains such as e-commerce platforms (Zhou et al., 2018; Singer et al., 2022), music recommendation services (Dai et al., 2024; Zhang et al., 2022), and video streaming applications (Lei et al., 2021; Zhang et al., 2024b). To accurately capture a user’s next interaction intent, sequential recommendation algorithms are designed to analyze historical interactions to mine underlying sequential patterns and model latent user preferences (Boka et al., 2024; Wang et al., 2019; Fang et al., 2020). Current mainstream sequential recommendation models, such as SASRec (Kang and McAuley, 2018) and UniSRec (Hou et al., 2022), adopt a Transformer-based architecture, leveraging their power attention mechanisms to adaptively weight past interacted items and use the final position’s encoded output as the user representation, as illustrated in Fig. 1(a). However, we argue this prevailing direct forward inference paradigm may lack nuanced comprehension of dynamic user preferences and evolving interest patterns, leading to suboptimal modeling for long-tail user interest and unpopular items.
@@
Recent studies from the natural language processing (NLP) community have demonstrated that Chain-of-Thought (CoT) during inference can significantly improve the performance of Large Language Models (LLMs) on complex tasks like mathematics and coding (Shao et al., 2024; Guo et al., 2024; Team et al., 2025; Wei et al., 2022). Specifically, when tackling intricate reasoning problems, these models engage in multi-step deep and deliberate logical analysis to generate the final response. This slow-thinking paradigm contrasts with traditional direct inference, mirroring human cognitive strategies before making decisions. Furthermore, Feng et al. (2023) theoretically uncover that the emergent thinking capabilities are attributed to the increased computational depth introduced by CoT reasoning, which allows models to overcome the expressivity limitations of direct answer even with constrained parameter sizes.
@@
Inspired by the above think-before-action paradigm, we introduce a novel reasoning-enhanced sequential recommendation framework, ReaRec, designed to empower SeqRec models with the capability to think before recommending. As illustrated in Fig. 1(b), our core idea is to perform a multi-step implicit thinking through reasoning chains before generating the final user representation, allowing the model to deepen the feature crossing in latent space. To prevent the recommender from confusing the sequence encoding stage and reasoning stage, we design a specialized positional encoding scheme to explicitly distinguish item representations from reasoning inputs. However, unlike NLP tasks, where explicit reasoning chains naturally provide process supervision to guide model optimization (Lightman et al., 2023; Luo et al., 2024; Setlur et al., 2024), implicit reasoning in sequential recommendation lacks effective intermediate signals. This absence of stepwise guidance could lead to unpredicted reasoning degradation issues, causing the recommender to either replicate prior reasoning patterns or progressively drift away from accurately modeling the user’s true interest distribution. Consequently, this may significantly impair the robustness and generalization capability of the recommendation model.
@@
Our extensive experiments on five benchmark datasets demonstrate the effectiveness of the proposed ReaRec framework. In particular, the ReaRec achieves an average performance gain of 7.49% across all metrics while incurring only 3.51% additional inference latency (cf. Sec. 4.2 and Sec. 4.3.3). Moreover, further analysis reveals several interesting empirical findings: (1) Enhancing modeling capability for underrepresented groups. The multi-step reasoning process steadily enhances the performance of users with sparse interactions and long-tail items. (2) Remarkable performance ceiling breakthrough. Post-hoc optimal reasoning step analysis shows that our framework elevates the performance ceilings for different backbone models by approximately 30%-50% (as shown in Fig. 2), highlighting its promising capability. We are optimistic that our proposed RecRec will open new avenues for exploring inference-time scaling for recommender systems.
@@ 2.2. Sequential Recommendation Pipeline
In a typical sequential recommendation pipeline, users’ historical interactions are first encoded into item embeddings. These item embeddings are then fed into a sequential model (e.g., transformer-based models) to produce a sequence representation, typically using the output from the final position (as illustrated in Fig. 1(a)). Finally, this sequence representation is used to calculate similarity scores with candidate item embeddings (such as dot product (Zhang et al., 2024a, b) or cosine similarity(Xu et al., 2025c; Li et al., 2023)) to predict interaction probabilities for the next item.
@@ 3. Methodology
In this section, we introduce ReaRec, a novel, simple, and highly scalable recommendation framework designed to unleash a model’s latent sequential reasoning capability. Instead of the traditional reasoning-free inference, our approach leverages multi-step implicit reasoning to refine user representations, fully exploiting the computational potential of sequential models to approximate the true distribution of user interests.
@@ 3.1.1. Self-attention Sequence Encoding
Given a user’s historical sequence $\mathcal{S}_{u}=[v^{u}_{1},v^{u}_{2},\ldots,v^{u}_{n}]$, we can obtain the item embeddings $\mathbf{h}_{u}\in\mathbb{R}^{n\times d}$ by looking up the embedding matrix $\mathbf{E}$. To fully leverage sequential information, we inject Absolute Position Embeddings into the item embeddings at the input layer. Specifically, for a given item $v$ at position $i$, the input representation is constructed by summing its item embedding $\mathbf{e}_{v}$ and the corresponding positional embedding $\mathbf{p}^{I}_{i}$:
@@ 3.1.2. Extended Inference-Time Reasoning
Specifically, rather than directly using $\mathbf{H}^{L}[-1]$ as the user representation, we autoregressively feed the hidden state of the last position back into the encoder for $\bm{K}$-pass forward computations. By effectively increasing (approximately $K$ times) inference-time computation, this approach further unleashes the model’s potential to capture intricate sequential dependencies. However, this inference strategy deviates from the original objective of sequential recommendation models, namely next-item prediction. To bridge this task gap, we introduce the Reasoning Position Embedding (RPE), denoted as $\mathbf{P}^{R}\in\mathbb{R}^{K\times d}$, to distinguish between the sequence encoding phase and the reasoning phase. At the $k$-th reasoning step, the model’s input embedding is defined as $\mathbf{H}^{0}\in\mathbb{R}^{(n+k-1)\times d}$. The first $n$ positions remain unchanged from the original input (i.e., Eq. (2)), while the latent representation $\mathbf{h}^{0}_{n+i}$ at position $n+i$ is calculated as the summation of the last output $\mathbf{h}^{L}_{n+i-1}$ from the previous step and the $i$-th reasoning position embedding $\mathbf{p}^{R}_{i}$:
@@
To differentiate between item encoding outputs and reasoning outputs, we denote the hidden states of the model’s final layer from position $n$ to $n+k$ as $\mathbf{R}=[\mathbf{r}_{0},\mathbf{r}_{1},\ldots,\mathbf{r}_{k}]$, where $\mathbf{r}_{i}\in\mathbb{R}^{d}$ represents the reasoning hidden state at the $i$-th step. To obtain the user representation, a straightforward approach is to follow the traditional paradigm and use the last reasoning output $\mathbf{r}_{K}$ as $\mathbf{h}_{u}$. Then, we calculate the predicted probability for the user $u$ as $\hat{y}=\text{softmax}(\mathbf{h}_{u}\cdot\mathbf{E}^{\top})$ and use cross-entropy loss as the recommendation objective function:
@@
However, this naive optimization objective still suffers from two critical issues: the lack of supervision signals for intermediate reasoning states makes the model susceptible to the risk of reasoning pattern degradation. Next, we introduce two simple yet effective reasoning learning strategies to address these challenges.
@@ 3.2. Ensemble Reasoning Learning (ERL)
To provide effective supervised signals for the implicit reasoning process, we propose an Ensemble Reasoning Learning (ERL) method. This approach uses the hidden states of different reasoning steps as multi-view representations of the user’s evolving interests. In other words, we apply the idea of ensemble learning (Dong et al., 2020; Sagi and Rokach, 2018) to aggregate diverse reasoning results from different steps, thereby avoiding suboptimal performance caused by the final output alone.
@@ 3.2.2. KL Divergence Regularization
(7) $\mathcal{L}_{\text{KL}}=\sum_{i=0}^{K-1}\sum_{j=i+1}^{K}\text{KL}(\hat{y}^{(i)}\|\hat{y}^{(j)}).$
@@ 3.3.1. Progressive Temperature Annealing (PTA)
Drawing an analogy the human cognitive process, as the thinking depth increases, reasoning pathways become progressively refined until converging toward optimal solutions. Similarly, we expect that as the number of recommendation model’s computations increases, the recommender would gradually clarify the user’s interest evolving patterns, which is manifested as sharper predicted distributions. Inspired by this motivation, we propose a simple Progressive Temperature Annealing (PTA) method to guide the reasoning process. To achieve this, we first introduce a temperature coefficient, $\tau_{k}$, for the $k$-th reasoning step to adjust the predicted distribution sharpness, which is formulated as follows:
@@ 3.3.2. Reasoning-aware Contrastive Learning (RCL)
where $H(\cdot)$ and $H(\cdot|\cdot)$ denote the entropy and conditional entropy of the random variable, respectively. By maximizing the MI between the original clean hidden states $\mathbf{R}$ and the denoised hidden states $\mathbf{\tilde{R}}$, it can effectively force the model to capture the essential sequential information from the user behavior data and historical reasoning process, achieving self-reflection in the implicit thought space.
@@ 4. Experiments
Table 1. The statistics of datasets.
@@ 4.3.1. Robustness Analysis Across User and Item Subgroups
To further analyze the robustness of our proposed ReaRec framework, we split users and items into different subgroups to gain deeper insights into the performance of the multi-step reasoning framework. Specifically, for users, we divide users into four equal-sized groups based on sequence length: {UG-0, UG-1, UG-2, UG-3}, where higher group numbers indicate longer sequences. For items, following previous work (Yang et al., 2023; Tang et al., 2024), we group them into four groups based on interaction frequency: {IG-0, IG-1, IG-2, IG-3}, where higher group numbers indicate more popular items. We ensure each item group contains the same sample numbers. The detailed experimental results are shown in Fig. 4.
We can clearly observe distinct performance trends across different user and item subgroups. For long-sequence user groups and unpopular item groups, recommendation quality (NDCG@20) tends to steadily improve as the reasoning steps increase. For example, in the item group IG-1, more reasoning steps bring better performance gains of 12.08%, 16.35%, and 18.69%, respectively. In contrast, performance tends to decline for users with long interaction sequences and popular items as the reasoning steps increase. We speculate that this is primarily because longer user sequences provide richer contextual information, making it easier to mine interest evolution patterns. Beyond a certain point, additional inference computation fails to yield further performance improvements and even leads to performance degradation due to overthinking. Similarly, for high-popularity items, their well-trained representations allow the recommender to easily capture collaborative signals, making deeper feature crossing depth less beneficial. Overall, long-tail users and items usually require more thinking space to reason sparse interaction signals, whereas highly active users and items may not need redundant computational expansion. In the future, it may be necessary to develop differentiated fast and slow reasoning mechanism for different user sequences to further improve overall recommendation performance.
@@ 4.3.2. Impact of Reasoning Steps on Recommendation Performance
We investigate the variation trend of recommendation performance under different inference steps, using NDCG@20 as the main evaluation metric. We compare the following approaches: (1) Base: The original SASRec sequential recommender serves as the baseline without reasoning enhancement; (2) Naive: Based on the Base method, we extend it to a multi-step reasoning paradigm, where the last hidden state is autoregressively fed back into the model, and only the final position is used directly as the user representation; (3) RPE: Building on the Naive approach, we further integrate Reasoning Positional Embeddings to bridge the task gap between sequence encoding mode and reasoning mode. Additionally, we also explore the performance of (4) Ensemble Reasoning Learning (ERL) and (5) Progressive Reasoning Learning (PRL) under multi-step reasoning.
@@ 5.2. Inference-time Reasoning
As the scaling law of large language models (LLMs) at the training stage has gradually reached its bottleneck (Chen et al., 2025; Snell et al., 2024; Kumar et al., 2025), researchers have shifted their focus toward inference-time scaling. Notably, OpenAI’s O1 series (Jaech et al., 2024), Qwen’s QwQ series (Team, 2024), and DeepSeek’s R1 series (Guo et al., 2025) have emerged as key milestone works, marking the transition from conversational AI to reasoning-intensive AI and opening promising pathways toward achieving Artificial General Intelligence (AGI). These works leverage emerging long Chain-of-Thought (CoT) mechanisms to reveal excellent test-time scaling phenomena—where increased computational power (via generating more tokens) during inference substantially improves the model’s problem-solving abilities (Wei et al., 2022; Besta et al., 2024; Yao et al., 2023; Yang et al., 2024; Wang et al., 2022; Chu et al., 2023). Extended reasoning space enables depth-scalable exploration, manifesting self-reflection capabilities (e.g., “Aha Moments”) that surpass traditional short-chain reasoning limitations (Chen et al., 2025; Ji et al., 2025). Compared with prior methods constrained by token-by-token generation in the discrete language spaces, which restricts the model’s expressive ability, another research direction focuses on implicit chain of thought reasoning in latent spaces, achieving both efficiency and performance gains in large language models (Hao et al., 2024; Xu et al., 2025a; Geiping et al., 2025; Chen et al., 2024) and multimodal foundation models (Shen et al., 2025; He et al., 2024). For example, Coconut (Hao et al., 2024) introduces continuous thinking in the latent reasoning space of LLMs, while Heima (Shen et al., 2025) compresses the entire multimodal CoT process into a single high-level thinking token (i.e., ¡CoT¿) to eliminate redundant intermediate token generation. Inspired by this think-before-action paradigm, we pioneer the exploration of implicit reasoning-enhanced sequential recommendation framework in this paper, proposing two lightweight reasoning learning methods to push the performance ceiling of sequential recommenders.

### v2

@@ 1. Introduction
Recommender systems (RS) have become ubiquitous in modern daily life, powering personalized services across domains such as e-commerce platforms (Zhou et al., 2018; Singer et al., 2022), music recommendation services (Dai et al., 2024; Zhang et al., 2022), and video streaming applications (Lei et al., 2021; Zhang et al., 2024b). To accurately capture a user’s next interaction intent, sequential recommendation algorithms are designed to analyze historical interactions to mine underlying sequential patterns and model latent user preferences (Boka et al., 2024; Wang et al., 2019; Fang et al., 2020). Current mainstream sequential recommendation models, such as SASRec (Kang and McAuley, 2018) and UniSRec (Hou et al., 2022), adopt a Transformer-based architecture, leveraging their power attention mechanisms to adaptively weight past interacted items and use the final position’s encoded output as the user representation, as illustrated in Fig. 1(a). However, we argue this prevailing direct forward inference paradigm may lack nuanced comprehension of dynamic user preferences and evolving interest patterns, leading to suboptimal modeling for long-tail user interest and unpopular items. Despite their efficiency, we argue that these direct inference paradigms often fall short in modeling long-tail users with fewer interactions and less popular items—scenarios that inherently demand more nuanced reasoning and deeper representation learning.
@@
Recently, many studies from the natural language processing (NLP) community have demonstrated that Chain-of-Thought (CoT) during inference can significantly improve the performance of Large Language Models (LLMs) on complex tasks like mathematics and coding (Shao et al., 2024; Guo et al., 2024; Team et al., 2025; Wei et al., 2022). By allowing the model to perform multi-step deliberation before generating a final output, CoT-based reasoning enhances the model’s capacity to handle complex problems beyond what direct inference allows. Furthermore, Feng et al. (2023) theoretically uncover that the emergent thinking capabilities are attributed to the increased computational depth introduced by CoT-based reasoning, which allows models to overcome the expressivity limitations of direct answer even with constrained parameter sizes.
@@
Motivated by these insights, we explore whether a similar think-before-action paradigm can benefit sequential recommendation, especially for challenging cases such as long-tail users and items. We propose ReaRec, a novel reasoning-enhanced framework that enables SeqRec models to engage in implicit multi-step reasoning during inference. As shown in Fig. 1(b), ReaRec performs autoregressive reasoning over latent representations before producing the final user embedding, thereby deepening feature crossing and improving representational richness. To prevent the recommender from confusing the sequence encoding stage and reasoning stage, we design a specialized positional encoding scheme to explicitly distinguish item representations from reasoning inputs. However, unlike NLP tasks, where explicit reasoning chains naturally provide process supervision to guide model optimization (Lightman et al., 2023; Luo et al., 2024; Setlur et al., 2024), implicit reasoning in sequential recommendation lacks effective intermediate signals. This absence of stepwise guidance could lead to unpredicted reasoning degradation issues, causing the recommender to either replicate prior reasoning patterns or progressively drift away from accurately modeling the user’s true interest distribution. Consequently, this may significantly impair the robustness and generalization capability of the recommendation model.
@@
Our extensive experiments on five benchmark datasets demonstrate the effectiveness of the proposed ReaRec framework. In particular, the ReaRec achieves an average performance gain of 7.49% across all metrics while incurring only 3.51% additional inference latency (cf. Sec. 4.2 and Sec. 4.3.3). Moreover, further analysis reveals several interesting empirical findings: (1) Enhancing modeling capability for underrepresented groups. The multi-step reasoning process steadily enhances the recommendation quality of users with sparse interactions and long-tail items. (2) Remarkable performance ceiling breakthrough. Post-hoc optimal reasoning step analysis shows that our method elevates the performance ceilings for different backbone models by approximately 30%-50% (as shown in Fig. 2), highlighting its promising capability. We are optimistic that our proposed RecRec will open new avenues for exploring inference-time scaling for recommender systems.
@@ 2.2. Sequential Recommendation Pipeline
In a typical sequential recommendation pipeline, users’ historical interactions are first encoded into item embeddings. These item embeddings are then fed into a sequential model (e.g., transformer-based models) to produce a sequence representation, typically using the output from the final position (as illustrated in Fig. 1(a)). Finally, this sequence representation is used to calculate similarity scores with candidate item embeddings (such as dot product (Zhang et al., 2024a, b) or cosine similarity(Xu et al., 2025c; Li et al., 2023)) to predict the probability of the user’s interaction with the next item.
@@ 3. Methodology
In this section, we introduce ReaRec, a novel, simple, and highly scalable recommendation framework designed to unleash a model’s latent sequential reasoning capability. Instead of the traditional direct recommendation without reasoning, our approach leverages multi-step implicit reasoning to refine user representations, fully exploiting the computational potential of sequential models to approximate the true distribution of user interests.
@@ 3.1.1. Self-attention Sequence Encoding
Given a user’s historical sequence $\mathcal{S}_{u}=[v^{u}_{1},v^{u}_{2},\ldots,v^{u}_{n}]$, we can obtain the item embeddings of these $n$ items by looking up the embedding matrix $\mathbf{E}$. To fully leverage sequential information, we inject Absolute Position Embeddings into the item embeddings at the input layer. Specifically, for a given item $v$ at position $i$, the input representation is constructed by summing its item embedding $\mathbf{e}_{v}$ and the corresponding positional embedding $\mathbf{p}^{I}_{i}$:
@@ 3.1.2. Extended Inference-Time Reasoning
Specifically, rather than directly using $\mathbf{H}^{L}[-1]$ as the user representation, we autoregressively feed the hidden state of the last position back into the encoder for $\bm{K}$-pass forward computations. By effectively increasing inference-time computation, this approach further unleashes the model’s potential to capture intricate sequential dependencies. However, this inference strategy deviates from the original objective of sequential recommendation models, namely next-item prediction. To bridge this task gap, we introduce the Reasoning Position Embedding (RPE), denoted as $\mathbf{P}^{R}\in\mathbb{R}^{K\times d}$, to distinguish between the sequence encoding phase and the reasoning phase. At the $k$-th reasoning step, the model’s input embedding is defined as $\mathbf{H}^{0}\in\mathbb{R}^{(n+k-1)\times d}$. The first $n$ positions remain unchanged from the original input (i.e., Eq. (2)), while the latent representation $\mathbf{h}^{0}_{n+i}$ at position $n+i$ is calculated as the summation of the last output $\mathbf{h}^{L}_{n+i-1}$ from the previous step and the $i$-th reasoning position embedding $\mathbf{p}^{R}_{i}$:
@@
To differentiate between item encoding outputs and reasoning outputs, we denote the hidden states of the model’s final layer from position $n$ to $n+k$ as $\mathbf{R}=[\mathbf{r}_{0},\mathbf{r}_{1},\ldots,\mathbf{r}_{k}]$, where $\mathbf{r}_{i}\in\mathbb{R}^{d}$ represents the reasoning hidden state at the $i$-th step. To obtain the user representation, a straightforward approach is to follow the traditional paradigm i.e., use the last reasoning output $\mathbf{r}_{K}$ as $\mathbf{h}_{u}$. Then, we calculate the predicted probability for the user $u$ as $\hat{y}=\text{softmax}(\mathbf{h}_{u}\cdot\mathbf{E}^{\top})$ and use cross-entropy loss as the recommendation objective function:
@@
However, this naive optimization objective still faces a critical issue: the lack of supervision signals for intermediate reasoning states makes the model vulnerable to the risk of reasoning pattern degradation. Next, we introduce two simple yet effective reasoning learning strategies to address these challenges.
@@ 3.2. Ensemble Reasoning Learning (ERL)
To provide effective supervised signals for the implicit reasoning process, we propose an Ensemble Reasoning Learning (ERL) method. This approach uses the hidden states of different reasoning steps as multi-view representations of the user’s evolving interests. In other words, we apply the idea of ensemble learning (Dong et al., 2020; Sagi and Rokach, 2018) to aggregate diverse reasoning results from different reasoning steps, thereby avoiding suboptimal performance caused by the final output alone.
@@ 3.2.2. KL Divergence Regularization
(7) $\mathcal{L}_{\text{KL}}=-\sum_{i=0}^{K-1}\sum_{j=i+1}^{K}\text{KL}(\hat{y}^{(i)}\|\hat{y}^{(j)}).$
@@ 3.3.1. Progressive Temperature Annealing (PTA)
Drawing an analogy the human cognitive process, as the thinking depth increases, reasoning pathways become progressively refined until converging toward optimal solutions. Similarly, we expect that as the model’s computations increases, the recommender would gradually clarify the user’s interest evolving patterns, which is manifested as sharper predicted distributions. Inspired by this motivation, we propose a simple Progressive Temperature Annealing (PTA) method to guide the reasoning process. To achieve this, we first introduce a temperature coefficient, $\tau_{k}$, for the $k$-th reasoning step to adjust the predicted distribution sharpness, which is formulated as follows:
@@ 3.3.2. Reasoning-aware Contrastive Learning (RCL)
where $H(\cdot)$ and $H(\cdot|\cdot)$ denote the entropy and conditional entropy of the random variable, respectively. By maximizing the MI between the original hidden states $\mathbf{R}$ and the denoised hidden states $\mathbf{\tilde{R}}$, it can effectively force the model to capture the essential sequential information from the user behavior data and historical reasoning process, achieving self-reflection in the implicit thought space.
@@ 4. Experiments
Table 1. The statistics of experimental datasets.
@@ 4.3.1. Robustness Analysis Across User and Item Subgroups
To further analyze the robustness of our proposed ReaRec framework, we split users and items into different subgroups to gain deeper insights into the performance of the multi-step reasoning framework. Specifically, for users, we divide users into four equal-sized groups based on sequence length: {UG-0, UG-1, UG-2, UG-3}, where higher group numbers indicate longer sequences. For items, following previous work (Yang et al., 2023; Tang et al., 2024), we group them into four groups based on interaction frequency: {IG-0, IG-1, IG-2, IG-3}, where higher group numbers indicate more popular items. We ensure each item group contains the same sample numbers. We fix the reasoning steps for PRL method during training at three, and analyze how recommendation performance changes for different user and item groups as reasoning steps increase during the inference phase. The detailed experimental results are shown in Fig. 4.
We can clearly observe distinct performance trends across different user and item subgroups. For short-sequence user groups and unpopular item groups, recommendation quality (NDCG@20) tends to steadily improve as the reasoning steps increase. For example, in the item group IG-1, more reasoning steps bring better performance gains of 12.08%, 16.35%, and 18.69%, respectively. In contrast, performance tends to decline for users with long interaction sequences and popular items as the reasoning steps increase. We speculate that this is primarily because longer user sequences provide richer contextual information, making it easier to mine interest evolution patterns. Beyond a certain point, additional inference computation fails to yield further performance improvements and even leads to performance degradation due to overthinking. Similarly, for high-popularity items, their well-trained representations allow the recommender to easily capture collaborative signals, making deeper feature crossing depth less beneficial. Overall, long-tail users and items usually require more thinking space to reason sparse interaction signals, whereas highly active users and items may not need redundant computational expansion. In the future, it may be necessary to develop differentiated fast and slow reasoning mechanism for different user sequences to further improve overall recommendation performance.
@@ 4.3.2. Impact of Reasoning Steps on Recommendation Performance
We investigate the variation trend of recommendation performance under different inference steps, that is, we train and perform inference using specified numbers of reasoning steps. We adopt NDCG@20 as the main evaluation metric. We compare the following approaches: (1) Base: The original SASRec sequential recommender serves as the baseline without reasoning enhancement; (2) Naive: Based on the Base method, we extend it to a multi-step reasoning paradigm, where the last hidden state is autoregressively fed back into the model, and only the final position is used directly as the user representation; (3) RPE: Building on the Naive approach, we further integrate Reasoning Positional Embeddings to bridge the task gap between sequence encoding mode and reasoning mode. Additionally, we also explore the performance of (4) Ensemble Reasoning Learning (ERL) and (5) Progressive Reasoning Learning (PRL) under multi-step reasoning.
@@ 5.2. Inference-time Reasoning
As the scaling law of large language models (LLMs) at the training stage has gradually reached its bottleneck (Chen et al., 2025; Snell et al., 2024; Kumar et al., 2025), researchers have shifted their focus toward inference-time scaling. Notably, OpenAI’s O1 series (Jaech et al., 2024), Qwen’s QwQ series (Team, 2024), and DeepSeek’s R1 series (Guo et al., 2025) have emerged as key milestone works, marking the transition from conversational AI to reasoning-intensive AI and opening promising pathways toward achieving Artificial General Intelligence (AGI). These works leverage emerging long Chain-of-Thought (CoT) mechanisms to reveal excellent test-time scaling phenomena—where increased computational power (via generating more tokens) during inference substantially improves the model’s problem-solving abilities (Wei et al., 2022; Besta et al., 2024; Yao et al., 2023; Yang et al., 2024; Wang et al., 2022; Chu et al., 2023; Zhang et al., 2025). Extended reasoning space enables depth-scalable exploration, manifesting self-reflection capabilities (e.g., “Aha Moments”) that surpass traditional short-chain reasoning limitations (Chen et al., 2025; Ji et al., 2025). Compared with prior methods constrained by token-by-token generation in the discrete language spaces, which restricts the model’s expressive ability, another research direction focuses on implicit chain of thought reasoning in latent spaces, achieving both efficiency and performance gains in large language models (Hao et al., 2024; Xu et al., 2025a; Geiping et al., 2025; Chen et al., 2024) and multimodal foundation models (Shen et al., 2025; He et al., 2024). For example, Coconut (Hao et al., 2024) introduces continuous thinking in the latent reasoning space of LLMs, while Heima (Shen et al., 2025) compresses the entire multimodal CoT process into a single high-level thinking token (i.e., ¡CoT¿) to eliminate redundant intermediate token generation. Inspired by this think-before-action paradigm, we pioneer the exploration of implicit reasoning-enhanced sequential recommendation framework in this paper, proposing two lightweight reasoning learning methods to push the performance ceiling of sequential recommenders.